OpenStudy (anonymous):

You roll a standard number cube. Find P(not 4).

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

there are six sides one is a four, five are not fours

OpenStudy (misty1212):

ratio of not fours to total

OpenStudy (anonymous):

\[ \Pr(\text{not}\ 4) = 1 - \Pr(4) \]

OpenStudy (misty1212):

lol

OpenStudy (anonymous):

these are my options 2/3 5/6 1/3 1/6

OpenStudy (misty1212):

number cube? what happened to the dice?

OpenStudy (er.mohd.amir):

getting 4 is p(4)=1/6 not p(4)=1-1/6 =5/6

OpenStudy (misty1212):

jeez it does not require subtractions five not fours to six \(\frac{5}{6}\)

OpenStudy (anonymous):

@misty1212 Please do not speak of die or dice, those that not hateful terms. They are called standard number cubes.